OpenStudy (anonymous):

kal1921 what is the perimeter of a rectangle if 2 sides are 5", and the other sides are both y? 5+y 10+2y 10y 20y

OpenStudy (solomonzelman):

\(\huge\color{magenta}{ \rm L }\) \(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid black ]{ \color{lightcyan}{\Huge\frac{~~~~~~~~\frac{~}{~~~\frac{~\frac{}{}~}{~\frac{~}{~}}~~~~}~~~}{~\frac{\frac{~}{~\frac{~\frac{}{~}}{~}}~~}{~}~~} } }}}\) \(\huge\color{red}{ \rm W }\) \(\huge\color{blue}{ \rm Perimeter = 2\times \left(\color{magenta}{ \rm L }+\color{red}{ \rm W } \right) }\) Or, you can use: \(\huge\color{blue}{ \rm Perimeter = 2\color{magenta}{ \rm L }+2\color{red}{ \rm W } }\)